# Victoria, BC, Canada

### v3

@@ 8 Conclusions
In this paper, we provide an overview of how mixed methods research designs can be used in software engineering, with the intention to inspire their use. Other resources, such as the ACM Sigsoft Empirical Standards sigsoft , as well as an earlier paper by Easterbrook et al. on selecting a research method in software engineering easterbrook_selecting_2008 , only lightly touch on the concepts and challenges of using mixed methods research in software engineering. As can be seen from Section 2, there is a long history of research behind the mixed method paradigm and use in other disciplines, some related to SE venkatesh2013bridging . Deciding what to include (and exclude) in this paper from this large body of research was challenging and we focused on the guiding principles, properties and designs that we hope will support software engineering researchers in applying mixed methods to lead to impactful insights. Still, applying mixed methods is not a trivial exercise, as we try to demonstrate through the identification of many possible antipatterns. There is a lot more to read and learn about mixed methods, and the references we provide too can provide a starting point for researchers to learn more.
The “selecting empirical research methods” paper easterbrook_selecting_2008 notably elaborates on topics such as theoretical stance and the role of theory in quantitative, qualitative and mixed research. The main takeaways from that paper also apply here: a mixed methods paper should have a clear research question that aligns with the research design; the methods used should answer this question; the limitations of any one method can be used to motivate the choice of other methods, but the limitations of and threats associated with mixing should also be reported; and the role of theory and whether the research is inductive or deductive should drive the research design and shape the integrated findings reported.
@@
We further describe the core properties that shape a diverse landscape of mixed methods research designs. These properties, which include the timing of methods, the inductive or deductive nature of the research, the research design planning or emergence, how data is integrated, and the mixed design, should be carefully considered up front and throughout the research process, and presented clearly in any research manuscripts. We do not provide a comprehensive description of all possible designs, but through several research scenarios, we show how the principles and properties can be considered as trade-offs and design choices to be made. The integration process throughout the research design is what sets mixed methods apart.

### v4

@@ 8 Conclusions
In this paper, we provide an overview of how mixed methods research designs can be used in software engineering, with the intention to inspire their use. Other resources, such as the ACM SIGSOFT Empirical Standards ralph2020empirical , as well as an earlier paper by Easterbrook et al. on selecting a research method in software engineering easterbrook_selecting_2008 , only lightly touch on the concepts and challenges of using mixed methods research in software engineering. As can be seen from Section 2, there is a long history of research behind the mixed method paradigm and use in other disciplines, some related to SE venkatesh2013bridging . Deciding what to include (and exclude) in this paper from this large body of research was challenging and we focused on the guiding principles, properties and designs that we hope will support software engineering researchers in applying mixed methods to lead to impactful insights. Still, applying mixed methods is not a trivial exercise, as we try to demonstrate through the identification of many possible antipatterns. There is a lot more to read and learn about mixed methods, and the references we provide too can provide a starting point for researchers to learn more.
The selecting empirical research methods paper by Easterbrook et al. easterbrook_selecting_2008 notably elaborates on topics such as theoretical stance and the role of theory in quantitative, qualitative and mixed research. The main takeaways from that paper also apply here: a mixed methods paper should have a clear research question that aligns with the research design; the methods used should answer this question; the limitations of any one method can be used to motivate the choice of other methods, but the limitations of and threats associated with mixing should also be reported; and the role of theory and whether the research is inductive or deductive should drive the research design and shape the integrated findings reported.
@@
We further describe the core properties that shape a diverse landscape of mixed methods research designs. These properties, which include the timing of methods, the inductive or deductive nature of the research, the research design planning or emergence, how data is integrated, and the mixed design, should be carefully considered up front and throughout the research process, and presented clearly in any research manuscript. We do not provide a comprehensive description of all possible designs, but through several research scenarios, we show how the principles and properties can be considered as trade-offs and design choices to be made. The integration process throughout the research design is what sets mixed methods apart.
@@
Acknowledgements
We are grateful to Prof. John Grundy for supporting a paper writing workshop at the HumaniSE Lab, Monash University, Melbourne in 2023. We also thank Cassandra Petrachenko and other members of the CHISEL Lab, UVic for feedback on the paper. This research was funded in part by the National Science and Engineering Research Council.